A rigid body rotates with 50 rad/sec about the axis π‘₯ = 𝑦 = 𝑧. Find the velocity and the acceleration of a point P of the solid. Application for P(1,1,0).


1
Expert's answer
2021-04-05T11:11:27-0400

"r=\frac{\sqrt{\begin{vmatrix}
 1& 1 \\
 1 & 1
\end{vmatrix}^2+\begin{vmatrix}
 1& 0 \\
 1 & 1
\end{vmatrix}^2+\begin{vmatrix}
 1& 0 \\
 1 & 1
\end{vmatrix}^2}}{\sqrt{1^2+1^2+1^2}}=\sqrt{\frac 23},"

"v=\omega r=50\cdot \sqrt{\frac 23}\approx 41~\frac ms,"

"a=\omega^2 r=50^2\cdot \sqrt{\frac 23}\approx 2000~\frac{m}{s^2}."